# references.bib

### v1

@@ Abstract
With the popularity of Hybrid Quantum-Classical architectures, particularly noisy intermediate-scale quantum (NISQ) architectures, comes the need for quality assurance methods tailored to their specific faults. In this study, we propose a taxonomy of faults in Hybrid Quantum-Classical architectures accompanied by a dataset of real faults in the identified categories. To achieve this, we empirically analysed open-source repositories for fixed faults. We analysed over 5000 closed issues on GitHub and pre-selected 529 of them based on rigorously defined inclusion criteria. We selected 133 faults that we labelled around symptoms and the origin of the faults. We cross-validated the classification and labels assigned to every fault between two of the authors. As a result, we introduced a taxonomy of real faults in Hybrid Quantum-Classical architectures. Subsequently, we validated the taxonomy through interviews conducted with eleven developers. The taxonomy was dynamically updated throughout the cross-validation and interview processes. The final version was validated and discussed through surveys conducted with an independent group of domain experts to ensure its relevance and to gain further insights.
1 Introduction
@@
Quantum Computing (QC) offers great promises for boosting current computational capabilities and enabling solutions to complex and challenging domain-specific problems [gheorghe-pop_quantum_2020]. In the last four decades, QC has evolved from theoretical concepts to real publicly available hardware and services. However, current computers are limited in their scalability to solve real-life calculations. While the main actors in realising the potential of QC have hitherto been physicists, computer scientists, and in particular, software engineers, are playing an increasing role in the following steps [miranskyy_testing_2019]. Current quantum hardware raises the need for specific software infrastructures and tools to mitigate its scalability issues [greiwe_effects_2023].
@@
For the foreseeable future, quantum systems are going to be interfaced and complemented with classical computers to achieve large-scale tasks. This has led to a wide-scale adoption of a new type of software architecture called Hybrid Quantum-Classical architectures. In particular, the current stage of quantum computing is termed the Noisy Intermediate-Scale Quantum (NISQ) era, where noisy quantum computation is wrapped around an architecture involving classical optimisation. Several algorithms have been developed to exploit the potential of current computers and are referred to as NISQ algorithms [lau_nisq_2022], which are but instances of Hybrid Quantum-Classical architectures. Until we have access to full-scale fault-tolerant quantum computing, we expect a quantum advantage for domain-specific NISQ algorithms [herrmann_quantum_2023]. However, even in the presence of fault-tolerant quantum computing, complex hybrid-quantum classical architectures are likely to remain prominent because of the complementary nature of these two computing paradigms. With such complex architectures comes the need for tailored quality assurance methods and tools.
@@
To respond to this need, in this study, we propose a fisrt taxonomy of real faults in Hybrid Quantum-Classical architectures. Although it is focusing on NISQ algorithms, we expect our taxonomy to apply to future Hybrid Quantum-Classical architectures, potentially including fault-tolerant quantum computing. We consider NISQ algorithms, which represent an early version of Hybrid Quantum-Classical architectures, to be an interesting case study to understand the behaviour of Quantum Systems when they are embedded in Classical Computations. To develop this taxonomy, we considered over 5000 Github issues, preselected 529 of them using rigorously defined inclusion criteria, and analysed them between two authors. We cross-validated our results and selected 133 real faults in Hybrid Quantum-Classical architectures. We then structured the results into a taxonomy and further labelled the faults in each classification with more specific descriptions of the nature of the fault. The initial version of the taxonomy was validated and continuously updated through interviews with eleven domain experts. The outcome of the interviews was validated once more through surveys involving a separate group of experts.
@@
    The integration of quantum circuits into a loop of classical calculations forms the structure of Hybrid Quantum-Classical architectures. It is natural to expect unique fault categories associated with it, pertaining to the quantum or classical components and their interfaces. We see similarities between such architectures and machine learning architectures, which motivated us to consider similar efforts in that domain [humbatova_taxonomy_2019] and analyse the similarities in the outcome of our research.
@@
Those questions motivated our analysis of real Hybrid Quantum-Classical system faults. The contributions of this paper are three-folded.
@@ 2 Background
2.1 Quantum Computing
Quantum computing (QC) was conceptualised to simulate quantum phenomena based on the postulates of quantum mechanics [feynman_simulating_1982]. It was later found to have potential applications that could offer significant speed-up over its classical counterpart. The resulting potential speed-up is often referred to as "quantum supremacy" [arute_quantum_2019]. However, demonstrating quantum supremacy on real hardware remains a long-standing challenge, especially at a scale where it would solve real-life calculations. Most agree this stage of QC will likely last for the next few years if not decades, and refer to it as the NISQ era [preskill_quantum_2018]. Regardless of the trajectory to large-scale fault-tolerant computing, quantum computers will have a different type of strength than classical computers and are going to be complemented by classical computers in hybrid architectures. NISQ algorithms are a prominent example that hybrid architecture combining small quantum circuits with classical computations could present some computational advantages [lau_nisq_2022]. Variational Quantum Algorithms (VQA) are the most common example of an efficient combination of a reduced quantum circuit inside a classical optimisation loop [tilly_variational_2022]. We present VQA as an example of Hybrid Quantum-Classical architecture in more detail at the end of this section.
2.2 Quantum Circuits
@@
A quantum circuit used in a Hybrid Quantum-Classical architecture typically begins with the initialisation of qubits and ends with their measurement, where the information is translated into classical bits. The purpose of a quantum circuit is to harness the principles of quantum mechanics, such as superposition and entanglement, to perform computational tasks that would be infeasible or significantly slower for “quantum parallelism”, i.e., “evaluating a function on many different values simultaneously” [nielsenChuang10].
Fig. 1 represents a simple quantum circuit, creating a three-qubit Greenberger-Horne-Zeilinger (GHZ) state, which is a specific entangled quantum state of three qubits. The GHZ state signifies that the three qubits are entangled in such a way that if the state of one qubit is measured, the states of the other two qubits are also determined. Input variables q[0], q[1], and q[3] represent the 3 input qubits. The horizontal line represents time, the blue elements are quantum logical gates, and the black elements are measurements where the qubit is observed and its state extracted into a classical bit. The maximum number of gates along all horizontal lines in a circuit is called depth. Fig. 1 represents a circuit of depth 3.
Figure 1: Simple Quantum circuit creating a GHZ state.
@@
There are different sources of noise in quantum computing, including unwanted entanglement among particles and imperfection in gates. Noise accumulates throughout the quantum circuit [clerk_introduction_2010]. Any quantum system suffers decoherence, induced by the interactions with its environment, causing it to lose its quantum behaviour over time [lau_nisq_2022]. Hence, current Hybrid Quantum-Classical architectures, e.g., NISQ algorithms, aim to reduce the number of qubits and the gate depth of the circuit to keep the noise level manageable.
2.3 Hybrid Quantum-Classical Architectures
@@
Hybrid Structure
Hybrid Quantum-Classical Architectures typically have four phases.
  1. 1.
    Classical pre-processing. The problem is translated into a cost function, and qubits are initialised.
  2. 2.
    Quantum Circuit. This is the core of a Hybrid Architecture. It typically has a low gate depth and few qubits to manage the noise level. Current Hybrid Architectures utilise around 50 qubits, and 1000 gates at maximum, with an average gate depth of 20 [lau_nisq_2022].
  3. 3.
    Loop of classical computation. The Quantum Circuit results are processed and a new quantum circuit is initialised. This process is repeated in a loop.
  4. 4.
    Classical post-processing. Once the desired results are reached, it is extracted and processed to answer the original problem.
@@
VQAs are among the most promising examples of NISQ algorithms [tilly_variational_2022]. The main goal of a VQA is to find the optimal parameters for a parameterized quantum circuit, leading to a solution for a given computational problem. We provide an overview of the structure and functioning of a VQA. This structure is typical of Hybrid Architecture. Future Fault-Tolerant Quantum Computing could replace current noisy hardware.
@@
    Optimisation Process. A parameterized quantum circuit, known as the variational ansatz, is designed. This circuit contains gates with adjustable parameters, denoted as $\theta$. It is initialised with an input $\theta_{0}$. A quantum state |$\ket{\psi(\theta)}$ is measured, and the outcomes are used to compute the expectation value of the objective function $O(\theta,\{\braket{H}_{U(\theta)}\})$.
@@
An in-depth explanation and a figure detailing figure can be found in the work of Bharti et al. [bharti_noisy_2022]. Fig. 2 [tilly_variational_2022] is a diagrammatic representation of a VQA in which each of the described steps can be found.
Figure 2: Detailed structure of a VQA [tilly_variational_2022]
3 Related work
3.1 Classical Computing
@@
Building datasets of reproducible real faults is a common and valuable practice in classical software engineering. These datasets offer a more realistic and eclectic base for software testing research on artificially generated faults. Noticeably, Defects4J gathers reproducible real faults in Java [just_defects4j_2014]. It was used in several studies, for instance, to inform the design of testing frameworks [martinez_automatic_2017] and to experiment with software engineering techniques [aleti_defects4j_2020]. We expect our study could similarly benefit research in Hybrid Quantum-Classical architectures and serve as an experimentation base for future studies.
@@
Taxonomies are a common way to propose a structured overview of a given problem in software engineering [UsmanEtAl017]. A study that inspired our work proposes a taxonomy of real faults in deep learning systems [humbatova_taxonomy_2019]. They analysed GitHub commits and issues for popular DL frameworks and related Stack Overflow posts to propose a taxonomy of real DL faults. They validated their work by interviewing 20 developers and integrating feedback received to their results. Our work follows a similar methodology applied to Hybrid Quantum-Classical architecture. Building a fault dataset and a taxonomy is a process that can be transferred to a different branch of software engineering since the data available online is similar. Also, the developers’ interview structure can be easily adapted to another speciality.
3.2 Quantum Computing
@@
Quantum software testing is challenging. Several studies point to a growing need for developing testing and debugging tools specific to quantum programs [miranskyy_testing_2019][paltenghi_bugs_2022][metwalli_tool_2022][gill_quantum_2022][pontolillo_multi-lingual_2023]. Some benchmarks of quantum bugs and bug fixes are available in the literature [zhao_bugs4q_2021][campos_qbugs_2021][luo_comprehensive_2022], as well as a low-level Quantum Benchmark Suite, QASMBench, focusing on NISQ Evaluation and Simulation [li_qasmbench_2023]. Zhao et al. [zhao_identifying_2021] identified some bug patterns in Quantum Programs after an in-depth analysis of Qiskit programs. In our study, we used one of them, namely, the Bugs4Q dataset [zhao_bugs4q_2021] to validate our search process. Although Bugs4Q has a broad scope and does not focus on Hybrid Quantum-Classical architectures, it contains one NISQ bug. In their systematic literature review, Gill et al. [gill_quantum_2022] briefly mention Hybrid Quantum-Classical architectures as a priority research area, without mentioning testing techniques for them. To our knowledge, real faults in Hybrid Quantum-Classical architectures, such as NISQ algorithms, have not been investigated yet. The closest study to ours concerns faults in Quantum Machine Learning, which is a special case of a NISQ algorithm [zhao_empirical_2023]. We used both datasets to validate our search query and process. We notice a growing interest in recent literature in Quantum testing and quality assurance, and very recently some studies have started to emphasise their investigation around NISQ and Hybrid Quantum-Classical architectures.

### v2

@@ Abstract.
With the popularity of Hybrid Quantum-Classical architectures, particularly noisy intermediate-scale quantum (NISQ) architectures in the gate-based quantum computing paradigm, comes the need for quality assurance methods tailored to their specific faults. In this study, we propose a taxonomy of faults in gate-based Hybrid Quantum-Classical architectures accompanied by a dataset of real faults in the identified categories. To achieve this, we empirically analysed open-source repositories for fixed faults. We analysed over 5000 closed issues on GitHub and pre-selected 529 of them based on rigorously defined inclusion criteria. We selected 133 faults that we labelled around symptoms and the origin of the faults. We cross-validated the classification and labels assigned to every fault between two of the authors. As a result, we introduced a taxonomy of real faults in gate-based Hybrid Quantum-Classical architectures. Subsequently, we validated the taxonomy through interviews conducted with eleven developers. The taxonomy was dynamically updated throughout the cross-validation and interview processes. The final version was validated and discussed through surveys conducted with an independent group of domain experts to ensure its relevance and to gain further insights.
Quantum Software Engineering, Quantum Software Testing, Fault Classification, Hybrid Quantum-Classical Systems, Taxonomy of Real Faults
$c$s: Software and its engineering, Computer systems organization
1. Introduction
@@
Quantum Computing (QC) holds the potential to enhance computational capabilities and address certain complex domain-specific problems that are intractable for classical methods (Gheorghe-Pop et al., 2020). In the last four decades, QC has evolved from theoretical concepts to real publicly available hardware and services. However, current computers are limited in their scalability to solve real-life calculations. While the main actors in realising the potential of QC have hitherto been physicists, computer scientists, and in particular, software engineers, are playing an increasing role in the following steps (Miranskyy and Zhang, 2019). Current quantum hardware raises the need for specific software infrastructures and tools to mitigate its scalability issues (Greiwe et al., 2023).
@@
For the foreseeable future, quantum systems—specifically gate-based quantum computers—are going to be interfaced and complemented with classical computers to achieve large-scale tasks. This has led to a wide-scale adoption of a new type of software architecture called Hybrid Quantum-Classical architectures. In particular, the current stage of quantum computing is termed the Noisy Intermediate-Scale Quantum (NISQ) era, where noisy quantum computation is wrapped around an architecture involving classical optimisation. Several algorithms have been developed to exploit the potential of current computers and are referred to as NISQ algorithms (Lau et al., 2022), which are instances of Hybrid Quantum-Classical architectures based on the gate-model quantum computing paradigm. Until we have access to full-scale fault-tolerant quantum computing, we expect a quantum advantage for domain-specific NISQ algorithms (Herrmann et al., 2023). However, even in the presence of fault-tolerant quantum computing, complex Hybrid-Quantum Classical architectures are likely to remain prominent because of the complementary nature of these two computing paradigms. The interest for Hybrid Quantum–Classical architectures has grown significantly over the past decade, as demonstrated by the number of related publications depicted in Figure 1. Figure 1a displays the yearly total number of publications resulting from a search query of “Hybrid Quantum Classical” keywords using arXiv’s tool from 2010 until today, while Figure 1b shows the same trend in peer-reviewed venues (Springer, ACM). We observe a clear growth with a significant increasing trend, reflecting the rising interest and investment in hybrid architectures as a promising use case of quantum computing. This trend likely stems from a broader recognition that quantum computing will, in the near future, be embedded within existing classical infrastructures. As the field matures, hybrid software will play a crucial role in enabling the seamless integration of this new computational paradigm into current systems. Such growing interest highlights the need for tailored quality-assurance methods and tools. Indeed, Ramalho, de Souza & Chaim note “gaps … related to … input states with hybrid interfaces” and call for “testing and debugging techniques that take advantage of the unique quantum computing characteristics” (Ramalho et al., 2024). Similarly, Murillo et al. (Murillo et al., 2025) observe that specific testing methods are crucial for Quantum and Hybrid Systems to scale. . These insights underscore why hybrid-specific testing, debugging, and verification strategies are essential to understand and classify the common fault types in hybrid architectures — a central motivation for the taxonomy presented in the following section.
(a) Publication Trend in ArXiv from 2010 to today.
(b) Publication Trend in SPRINGER (in blue) and ACM (in orange) from 2010 to today.
Figure 1. Comparaison of Publication Trends for Hybrid Quantum Classical in ArXiv and Peer-reviewed venues from 2010 to today.
@@
To respond to this need, in this study, we propose a first taxonomy of real faults in Hybrid Quantum-Classical architectures. Although it is focusing on NISQ algorithms running on gate-based quantum hardware, we expect our taxonomy to apply to future Hybrid Quantum-Classical architectures, potentially including fault-tolerant quantum computing. We consider NISQ algorithms, which represent an early version of Hybrid Quantum-Classical architectures, to be an interesting case study to understand the behaviour of Quantum Systems when they are embedded in Classical Computations. To develop this taxonomy, we considered over 5000 Github issues, preselected 529 of them using rigorously defined inclusion criteria, and analysed them between two authors. We cross-validated our results and selected 133 real faults in Hybrid Quantum-Classical architectures. We then structured the results into a taxonomy and further labelled the faults in each classification with more specific descriptions of the nature of the fault. The initial version of the taxonomy was validated and continuously updated through interviews with eleven domain experts. The outcome of the interviews was validated once more through surveys involving a separate group of experts.
@@
    The integration of quantum circuits into a loop of classical calculations forms the structure of Hybrid Quantum-Classical architectures. It is natural to expect unique fault categories associated with it, pertaining to the quantum or classical components and their interfaces. We see similarities between such architectures and machine learning architectures, which motivated us to consider similar efforts in that domain (Humbatova et al., 2019) and analyse the similarities in the outcome of our research.
@@
Those questions motivated our analysis of real Hybrid Quantum-Classical architectures faults. The contributions of this paper are three-folded.
@@ 2. Background
2.1. Quantum Computing
Quantum computing (QC) was conceptualised to simulate quantum phenomena based on the postulates of quantum mechanics (Feynman, 1982). It was later found to have potential applications that could offer significant speed-up over its classical counterpart. The resulting potential speed-up is often referred to as ”quantum supremacy” (Arute et al., 2019). However, demonstrating quantum supremacy on real hardware remains a long-standing challenge, especially at a scale where it would solve real-life calculations. Most agree this stage of QC will likely last for the next few years if not decades, and refer to it as the NISQ era (Preskill, 2018). Regardless of the trajectory to large-scale fault-tolerant computing, quantum computers will have a different type of strength than classical computers and are going to be complemented by classical computers in Hybrid architectures. NISQ algorithms are a prominent example that Hybrid architecture combining small quantum circuits with classical computations could present some computational advantages (Lau et al., 2022). Variational Quantum Algorithms (VQA) are the most common example of an efficient combination of a reduced quantum circuit inside a classical optimisation loop (Tilly et al., 2022). We present VQA as an example of Hybrid Quantum-Classical architecture in more detail at the end of this section.
Different paradigms
In the landscape of quantum computing paradigms, gate-based quantum computing stands out as the most widely adopted and a general-purpose model. It relies on the manipulation of qubits through a sequence of quantum gates, analogous to logic gates in classical computing, and forms the foundation for most current quantum algorithms, including hybrid approaches such as Variational Quantum Algorithms (VQAs). In contrast, quantum annealing is a specialized model primarily designed for solving combinatorial optimization problems by leveraging quantum fluctuations to explore low-energy states of a system. Commercial implementations, such as those from D-Wave Systems, are based on this paradigm but are not universal and are limited in their applicability beyond specific problem classes. Another alternative is measurement-based quantum computing (MBQC), which executes computations by performing a series of measurements on a highly entangled initial resource state (typically a cluster state), rather than applying gates directly. While MBQC provides a theoretically equivalent model to gate-based systems, it is less mature in terms of hardware support and algorithmic development. Given the broader applicability and better compatibility with near-term hardware, gate-based quantum computing remains the dominant framework, particularly for NISQ-era hybrid algorithms.
2.2. Quantum Circuits
@@
A quantum circuit used in a Hybrid Quantum-Classical architecture typically begins with the initialisation of qubits and ends with their measurement, where the information is translated into classical bits. The purpose of a quantum circuit is to harness the principles of quantum mechanics, such as superposition and entanglement, to perform computational tasks that would be infeasible or significantly slower for “quantum parallelism”, i.e., “evaluating a function on many different values simultaneously” (Nielsen and Chuang, 2010).
Quantum Gates
Quantum circuits are constructed using quantum gates, which are unitary operations acting on one or more qubits. These gates manipulate the quantum state by exploiting quantum phenomena. Common single-qubit gates include the Hadamard (H) gate, which creates superposition, the Pauli-X gate, which acts like a classical NOT, and phase-shift gates such as S and T. Multi-qubit gates, such as the Controlled-NOT (CNOT) gate, are essential for creating entanglement between qubits. Together, these gates form a universal set capable of implementing any quantum computation when combined in appropriate sequences.
Example of a quantum circuit
Fig. 2 represents a simple quantum circuit, creating a three-qubit Greenberger-Horne-Zeilinger (GHZ) state, which is a specific entangled quantum state of three qubits. The GHZ state signifies that the three qubits are entangled in such a way that if the state of one qubit is measured, the states of the other two qubits are also determined. Input variables $q[0]$, $q[1]$, and $q[2]$ represent the 3 input qubits. The horizontal line represents time, the square and round elements are quantum logic gates, and the final boxes and arrow elements represent measurements, where the qubit is observed and its state is extracted into classical bits. The maximum number of gates along all horizontal lines in a circuit is called depth. Fig. 2 represents a circuit of depth 3.
Figure 2. Simple Quantum circuit creating a GHZ state.
@@
There are different sources of noise in quantum computing, including unwanted entanglement among particles and imperfection in gates. Noise accumulates throughout the quantum circuit (Clerk et al., 2010). Any quantum system suffers decoherence, induced by the interactions with its environment, causing it to lose its quantum behaviour over time (Lau et al., 2022). Hence, current Hybrid Quantum-Classical architectures, e.g., NISQ algorithms, aim to reduce the number of qubits and the gate depth of the circuit to keep the noise level manageable.
2.3. Hybrid Quantum-Classical Architectures
@@
Structure of Hybrid Quantum-Classical Architectures
The general workflow consists of four main phases:
  1. (1)
    Classical pre-processing. The problem of interest (e.g., optimisation, simulation, or classification) is mapped into a quantum-compatible representation. This typically involves formulating a cost function and encoding it into a parameterised quantum circuit, known as an Ansatz. Initial parameters are chosen—often heuristically or randomly—and qubits are prepared in a standard reference state such as $\left|0\right\rangle^{\otimes n}$.
  2. (2)
    Quantum circuit execution. A parameterised quantum circuit is executed with the current parameter values. The circuit consists of layers of single-qubit rotations and entangling gates and is typically kept shallow to mitigate noise. This is the core of a Hybrid Architecture. It typically has a low gate depth and few qubits to manage the noise level. Current Hybrid architectures utilise around 50 qubits, and 1000 gates at maximum, with an average gate depth of 20 (Lau et al., 2022). The circuit is run multiple times (shots) to generate measurement statistics.
  3. (3)
    Classical optimisation loop. Measurement outcomes are used to estimate the value of the cost function, which is then fed into a classical optimiser that updates the parameters. This process iterates, forming a hybrid feedback loop until a convergence criterion is met.
  4. (4)
    Classical post-processing. Once a sufficiently good solution is found, the output—such as optimised parameters or low-energy states—is further processed or interpreted using classical resources, depending on the application.
@@
VQAs are among the most promising examples of NISQ algorithms (Tilly et al., 2022). The main goal of a VQA is to find the optimal parameters for a parameterized quantum circuit, leading to a solution for a given computational problem. We provide an overview of the structure and functioning of a VQA. This structure is typical of Hybrid Architecture. Future Fault-Tolerant Quantum Computing could replace current noisy hardware.
@@
[…]
@@
An in-depth explanation and a detailing figure can be found in the work of Bharti et al. (Bharti et al., 2022). Fig. 3 (Tilly et al., 2022) is a diagrammatic representation of a VQA in which each of the described steps can be found.
Figure 3. Detailed structure of a VQA (Bharti et al., 2022)
3. Related work
This section positions our study within the broader context of software engineering and quantum computing research. We begin by reviewing established practices in classical software engineering, including the construction of real fault benchmarks and the development of taxonomies for fault classification. We then examine existing efforts in quantum software testing, with a particular focus on real faults in quantum and Hybrid Quantum-Classical architectures. Finally, we compare our study to the most closely related work in this area. A summary of these contributions and their limitations is presented in Table 1, highlighting the novelty and gap addressed by our work.
3.1. Classical Computing
@@
Building datasets of reproducible real faults is a common and valuable practice in classical software engineering. These datasets offer a more realistic and eclectic base for software testing research on artificially generated faults. Noticeably, Defects4J gathers reproducible real faults in Java (Just et al., 2014). It was used in several studies, for instance, to inform the design of testing frameworks (Martinez et al., 2017) and to experiment with software engineering techniques (Gay and Just, 2020). We expect our study could similarly benefit research in Hybrid Quantum-Classical architectures and serve as an experimentation base for future studies.
Use of Taxonomy in Software Engineering
In science and engineering, systematically organising and classifying knowledge advances the field by providing a common discourse; it facilitates knowledge sharing, clarifies relationships between concepts, identifies knowledge gaps, and supports decision-making. This classification helps researchers and practitioners generalise, communicate, and apply findings effectively. As stated in Usman et al.’s systematic mapping study of Taxonomies in Software Engineering (SE) (Usman et al., 2017a), SE being a broad and diverse field, it relies heavily on taxonomies structured classification schemes that describe terms and their relationships to organise its extensive knowledge base. In summary, taxonomies contribute to both theoretical clarity and practical applicability in the software engineering discipline.
@@
Taxonomies being a common way to propose a structured overview of a given problem in software engineering (Usman et al., 2017b), a study that inspired our work proposes a taxonomy of real faults in deep learning systems (Humbatova et al., 2019). They analysed GitHub commits and issues for popular DL frameworks and related Stack Overflow posts to propose a taxonomy of real DL faults. They validated their work by interviewing 20 developers and integrating feedback received to their results. Our work follows a similar methodology applied to Hybrid Quantum-Classical architecture. Building a fault dataset and a taxonomy is a process that can be transferred to a different branch of software engineering since the data available online is similar. Also, the developers’ interview structure can be easily adapted to another specialty.
3.2. Quantum Computing
@@
Quantum software testing is challenging. Several studies point to a growing need for developing testing and debugging tools specific to quantum programs (Miranskyy and Zhang, 2019)(Paltenghi and Pradel, 2022)(Metwalli and Van Meter, 2022)(Gill et al., 2022)(Pontolillo and Mousavi, 2023). Some benchmarks of quantum bugs and bug fixes are available in the literature (Zhao et al., 2021b)(Campos and Souto, 2021)(Luo et al., 2022), as well as a low-level Quantum Benchmark Suite, QASMBench, focusing on NISQ Evaluation and Simulation (Li et al., 2023). Zhao et al. (Zhao et al., 2021a) identified some bug patterns in Quantum Programs after an in-depth analysis of Qiskit programs. In our study, we used one of them, namely, the Bugs4Q dataset (Zhao et al., 2021b) to validate our search process. Although Bugs4Q has a broad scope and does not focus on Hybrid Quantum-Classical architectures, it contains one NISQ bug. In their systematic literature review, Gill et al. (Gill et al., 2022) briefly mention Hybrid Quantum-Classical architectures as a priority research area, without mentioning testing techniques for them. To our knowledge, real faults in Hybrid Quantum-Classical architectures, such as NISQ algorithms, have not been investigated yet. The closest study to ours concerns faults in Quantum Machine Learning, which is a special case of a NISQ algorithm (Zhao et al., 2023). We used both datasets to validate our search query and process. We notice a growing interest in recent literature in Quantum testing and quality assurance, and very recently some studies have started to emphasise their investigation around NISQ and Hybrid Quantum-Classical architectures.